### v1

# https://hertie.ai/

Abstract
In many medical imaging tasks, convolutional neural networks (CNNs) efficiently extract local features hierarchically. More recently, vision transformers (ViTs) have gained popularity, using self-attention mechanisms to capture global dependencies, but lacking the inherent spatial localization of convolutions. Therefore, hybrid models combining CNNs and ViTs have been developed to combine the strengths of both architectures. However, such hybrid CNN-ViT models are difficult to interpret, which hinders their application in medical imaging. In this work, we introduce an interpretable-by-design hybrid fully convolutional CNN-Transformer architecture for medical image classification. Unlike widely used post-hoc saliency methods for ViTs, our approach generates faithful and localized evidence maps that directly reflect the model’s decision process. We evaluated our method on two medical image classification tasks using color fundus images. Our model not only achieves state-of-the-art predictive performance compared to both black-box and interpretable models but also provides class-specific sparse evidence maps in a single forward pass.
Keywords:
Interpretability Vision Transformers Convolutional Neural Networks Hybrid architecture Dual Resolution Self-Attention.
1 Introduction
Convolutional neural networks (CNNs) are at the heart of many successful applications in medical image analysis [9], but more recently, vision transformers (ViTs) have emerged as a competitive alternative [10], demonstrating strong performance in medical imaging tasks [3, 22]. Although CNNs are highly effective at capturing complex local patterns in images, the size of their receptive field is smaller than some disease-related lesions [14]. In contrast, vision transformers leverage self-attention (SA) [23] to capture long-range dependencies, providing a more global understanding of the image. Despite these advantages, ViTs require substantial computational resources, often demanding large-scale datasets for effective training [16, 22], while also facing challenges in interpretability [13].
To address the weaknesses of both approaches, a promising alternative are hybrid CNN-Transformer architectures. Several studies have used such architectures [12, 15, 16, 22], improving performance for tasks that require combining local features with global relationships for classification. Yet, the interpretability of such hybrid approaches has remained an issue, as they require methods suitable for transformer architectures. To this end, either CNN-based methods have been adapted to ViTs [19, 4] or ViT-specific techniques have been proposed [1, 7, 8]. The most commonly used ViT-specific approach has been to visualize attention maps across layers, as these capture interactions between input regions. However, attention is not class-specific and merely illustrates relationships between input patches rather than their direct contribution to the model prediction [5, 20, 13]. Alternatively, post-hoc CNN-based methods like LRP [4] and GradCAM [19] have been successfully adapted to ViT by integrating gradients within the self-attention layers, offering class-wise explanations [8]. Yet, these are model-specific and struggle with hierarchical architectures like the Swin Transformer [17].
Here, we propose a novel inherently interpretable-by-design hybrid CNN-Transformer architecture for image classification that combines the feature extraction strengths of CNNs with the ability of ViTs to capture long-range dependencies from dual-resolution features. It integrates recent advances, including convolutional ViTs [26], dual-resolution self-attention [12], and sparse explanations [14]. We evaluated our model with different convolutional architectures as backbone (ResNet vs. BagNet) on two clinically relevant tasks using publicly available fundus image datasets for Diabetic Retinopathy (DR) detection and Age-Related Macular Degeneration (AMD) severity classification. Our model maintained high predictive performance compared to both interpretable and non-interpretable state-of-the-art models while providing faithful, interpretable explanations that accurately localize disease-related lesions, even under distribution shift, outperforming classical post-hoc methods.
2 Developing a self-explainable hybrid CNN-ViT model
Figure 1: Interpretable-by-design hybrid CNN-Transformer model. (a) Input image. The black patches illustrate the small receptive field of (b) the BagNet backbone. (c) High- and downsampled low-resolution feature maps are processed by a window SA module and then fused. (d) Final attention map. (e) Convolutional classifier applies $C$ kernels to generate (f) the class evidence map $\mathbf{A}$. (g) Predictions are obtained by averaging class evidence maps, while explanations are directly derived from (f).
2.1 Hybrid CNN-ViT architecture
In our hybrid architecture (Fig. 1), CNN and ViT modules were used sequentially, with the output of the CNN module directly serving as the input to the transformer module. Specifically, the CNN module acted as a feature extractor, capturing local patterns, while the ViT module modeled long-range dependencies between the extracted features, enhancing the model’s ability to understand broader contexts. Given an input image $\textbf{X}\in\mathbb{R}^{H\times W\times C}$ (with height $H$, width $W$, and the number of channels $C$), the CNN backbone (Fig. 1b) extracted a spatial feature representation $\mathbf{Z}=f_{\theta}(\mathbf{X})\in\mathbb{R}^{M\times N\times D}$, where $\theta$ denoted the model parameter, $M\times N$ is the spatial size, and $D$ as the feature dimension. We used a ResNet50 (receptive field: 427×427) or a BagNet-33 (33×33) as backbone. Unlike the ResNet, the BagNet aggregated only local features in a bag-of-words manner [6]. The transformer module (Fig. 1c) used a convolutional window self-attention (Conv-wSA) mechanism, which operated on both high- and low-resolution versions of the original feature maps to produce an attention map $\mathbf{W}=g_{\phi}(\mathbf{Z}_{h},\mathbf{Z}_{l})\in\mathbb{R}^{M\times N\times D}$. Here, $\mathbf{Z}_{h}=\mathbf{Z}$ denoted the high-resolution feature map, while $\mathbf{Z}_{l}=d(\mathbf{Z},r)\in\mathbb{R}^{\frac{M}{r}\times\frac{N}{r}\times D}$ was the low-resolution counterpart, obtained via the downsampling function $d(.)$ with a reduction factor $r$. The self-attention preserved the original high-resolution feature map size (Fig. 1d). The classification module (Fig. 1e) consisted of a convolutional layer with $C$ convolution kernels of size $1\times 1$, and unit stride, producing an evidence map $\mathbf{A}=h_{\psi}(\mathbf{W})\in\mathbb{R}^{M\times N\times C}$, where $C$ represents the number of classes and $\psi$ denotes the parameter of $h$. The final prediction was computed by applying spatial average pooling followed by a softmax operation: $\mathbf{\hat{y}}=\mbox{Softmax}\big{(}\mbox{AvgPool}(\mathbf{A})\big{)}\in\mathbb{R}^{1\times C}$. The result as a $C$-dimensional probability distribution representing class likelihoods.
2.2 Learning long-range dependencies with convolutional DRSA
To learn long-range dependencies between the convolutional features, we used a transformer module with dual resolution self-attention [12], for which the linear fully connected layer had been replaced by a convolutional layer [26] as follows: $\text{SA}_{h}=\text{Softmax}\big{(}\frac{\mathbf{Q}_{h}\mathbf{K^{T}_{h}}}{\alpha}\big{)}\mathbf{V}_{h}$, $\text{SA}_{l}=\text{Softmax}\big{(}\frac{\mathbf{Q}_{l}\mathbf{K^{T}_{l}}}{\alpha}\big{)}\mathbf{V}_{l}$ where $\alpha$ is the scaling factor, and $\mathbf{Q}_{h},\mathbf{K}_{h},\mathbf{V}_{h}$ and $\mathbf{Q}_{l},\mathbf{K}_{l},\mathbf{V}_{l}$ are the queries, keys, and value embeddings generated for $\mathbf{Z}_{h}$ and $\mathbf{Z}_{l}$ using convolutional operations. The final self-attention is computed as: $\text{SA}_{final}=\text{GDFN}_{\delta}\big{(}\text{Proj}_{\beta}(\text{SA}_{h}+\text{Up}(\text{SA}_{l}))\big{)}$ where $\text{Up}(\text{SA}_{l})$ upsamples $\text{SA}_{l}$, which was then aggregated with $\text{SA}_{h}$ and passed through a convolutional projection with parameter $\beta$. The representation was refined using a Gated-Dconv Feed-Forward Network (GDFN) [28] with parameter $\delta$, which enhanced spatial structures while suppressing irrelevant features, ensuring that only relevant information contributed to predictions and improving generalization.
2.3 Enhancing interpretability with a sparse convolutional classifier
In standard ViT and hybrid CNN-Transformer models, the classification head includes a fully connected layer (FCL), which discards spatial information, limiting interpretability. Our architecture addressed this by preserving spatial information using convolutional operations in the self-attention module, generating attention maps that captured long-range dependencies between regions in the same window. To enhance interpretability, we replaced the FCL with a convolutional classifier, referred to as the class evidence layer. This layer leveraged spatial information to produce class-wise evidence maps (Fig. 1f), where each pixel reflected the local contribution of input regions to the final prediction. After prediction, evidence maps were upsampled and overlaid on the input for visualization (Fig. 1g).
Furthermore, introducing an explicit class evidence layer allowed us to apply an $\ell_{1}$ sparsity constraint on the class evidence maps $\mathbf{A}_{c}$, enhancing interpretability [14]. This lead to the following loss function:
$\mathcal{L}(\mathbf{y},\mathbf{\hat{y}})=\mbox{CE}(\mathbf{y},\mathbf{\hat{y}})+\lambda\sum_{i,j,c}|\mathbf{A}_{c}^{ij}|.$ (1)
Here, CE denoted the cross-entropy loss, and $\mathbf{y}$ represented the reference class labels. The sparsity of the evidence maps was controlled by the hyperparameter $\lambda$. The entire model was trained end-to-end using gradient descent.
3 Results
3.1 Datasets
We used two publicly available retinal fundus datasets, the Kaggle Diabetic Retinopathy (DR) [11] and the Age-Related Eye Disease Study (AREDS) [21]. The Kaggle DR dataset had 45,923 images from 28,984 subjects after quality filtering with class distributions: 73% No DR, 15% Mild, 8% Moderate, 3% Severe, and 1% Proliferative DR. The AREDS dataset contained 34,079 images from 4,757 participants. AMD severity was grouped into six categories [2, 21]: 49%, 19% 14%, 3%, 12%, 1% for early, moderate, adv. intermediate, early late, active neovascular and end-stage AMD.
Images were resized to 512 × 512, normalized, and augmented with cropping, flipping, color jitter, and rotation. Datasets were split into 75% training, 10% validation, and 15% test, keeping each participant’s records in the same split. To evaluate our model’s ability to localize DR-related lesions, we used the IDRiD dataset [18], which includes 81 fundus images with pixel-level annotations for microaneurysms (MA), hemorrhages (HE), soft exudates (SE), and hard exudates (EX), offering insights into interpretability through localization performance.
Table 1: Classification performance on multiclass detection on the test sets.
AREDS AMD Kaggle Fundus DR
Parameters Accuracy $\kappa$ Accuracy $\kappa$
ViT $86,094,341$ $0.763$ $0.900$ $0.811$ $0.708$
Swin $86,883,709$ $0.780$ $\mathbf{0.914}$ $0.844$ $0.779$
ResNet $23,518,277$ $0.782$ $0.899$ $0.857$ $0.815$
BagNet $16,271,429$ $0.745$ $0.882$ $0.859$ $0.826$
ResNet-FCL-SA $136,845,927$ $0.780$ $0.900$ $0.859$ $0.822$
BagNet-FCL-SA $129,614,343$ $0.770$ $0.897$ $0.854$ $0.828$
ResNet-Conv-SA $69,737,863$ $0.786$ $0.906$ $0.854$ $0.830$
BagNet-Conv-SA $62,915,899$ $0.768$ $0.895$ $\mathbf{0.860}$ $\mathbf{0.838}$
sResNet-Conv-SA $69,737,863$ $\mathbf{0.786}$ $0.902$ $0.847$ $0.800$
sBagNet-Conv-SA $62,915,899$ $0.773$ $0.905$ $0.853$ $0.809$
3.2 Self-explanaible hybrid models achieved SOTA performance
Figure 2: Examples explanations. From left to right, heatmaps for the correctly predicted class. The first row shows an example (grade 1) from the Kaggle dataset, while the second row shows an example (grade 2) from the AREDS dataset.
We first evaluated our models on multiclass DR detection and AMD severity classification. Using a ResNet50 or a BagNet-33 as backbone, our model integrated a dual-resolution convolutional self-attention (DR-Conv-SA) and a GDFN module [28]. The reduction factor was set to $r=2$ using max pooling for improved performance. We tuned the window size ($w=10$ for BagNet, $w=8$ for ResNet) and the regularization coefficient $\lambda$ (Eq. 1) to balance the accuracy and sparsity of the class activation map.
We compared our sparse models to their dense version ($\lambda=0$), a version with linear SA and an FCL classifier, and other models including a ResNet50, a BagNet33, a ViT32 (with input size 384), and a Swin Transformer (with input size 384, patch size 4, window size 12, [24]). All models were initialized with pre-trained weights and trained with the same setup [^1] Code available at https://anonymous.4open.science/r/Expl-CNN-Transformer/: data augmentation, cross-entropy loss, cosine learning rate schedule, and SGD optimizer (learning rate $10^{-4}$, momentum 0.9, weight decay $5\cdot 10^{-4}$) for 70 epochs with a mini-batch size of 8 on an NVIDIA A40 GPU, using PyTorch, with the best models selected based on validation accuracy.
Our interpretable-by-design hybrid CNN-transformer models achieved state-of-the-art performance on both tasks, with the dense model with BagNet backbone providing the best DR classification results, while the model with ResNet backbone performed best ($\kappa$) for AMD severity classification (Tab.1). Despite the sparsity penalty on the class activation map, the sparse models achieved competitive accuracy with only slightly lower $\kappa$. Interestingly, for AMD detection, $\kappa$ was higher than the accuracy, which is likely because misclassifications primarily occur between similar classes (Fig. 3d).
3.3 Sparsity constraints enhance class evidence maps
Figure 3: Quantitative evaluation of heatmaps and confusion matrices. (a-c) Precision evaluation on IDRiD dataset. (e,f) Sensitivity analysis of different heatmaps for DR detection and AMD severity classification. (d,g,h) Confusion matrices of different models for DR detection and AMD severity classification on the test sets.
We next compared evidence maps from our model to attribution maps generated with GradCAM [19] on the ViT baseline. As these were multiclass tasks, we only showed class evidence maps from the correctly predicted class. Our class evidence maps, obtained from the convolutional layer before average pooling, clearly highlighted image features relevant to the predicted class (Fig. 2). We noticed that GradCAM on ViT produced cluttered, hard-to-interpret heatmaps. In contrast, the hybrid ResNet-Transformer generated coarser heatmaps due to its large receptive field, while the hybrid BagNet-Transformer provided more localized explanations. The sparse models further refined this by producing sparser heatmaps, focusing decisions on smaller yet relevant retinal regions.
3.4 Evidence maps provide faithful and localized explanations
We quantitatively assessed quality of the explanations by evaluating their precision in identifying DR lesions [14]. Following the International Clinical Diabetic Retinopathy Scale [25], we evaluated three cases: (a) binary evaluation (Fig. 3a), averaging disease-class heatmaps and combining all lesion annotations; (b) severe DR (Fig. 3b), where MAs, HEs, and SEs were combined, and the precision was computed from the severe grade heatmap (c) proliferative DR (Fig. 3c), were all lesions were combined and precision was evaluated from the heatmap from the proliferative grade heatmap. Precision was measured as the proportion of positively activated regions containing lesions [14], using $33\times 3$ 33 non-overlapping patches to match BagNet’s receptive field. For ViT and hybrid FCL models, GradCAM-generated heatmaps were used, and patches were extracted from positively activated regions.
In all cases, the sparse BagNet-Transformer showed considerably higher precision than all other models and outperformed the base BagNet, suggesting that incorporating attention improved both classification and interpretability. The ResNet-Transformer with an explicit class-evidence layer performed worse, likely due to its larger receptive field producing coarser localizations (Fig. 2).
Subsequently, we additionally measured the faithfulness of the explanations by evaluating their ability to identify relevant regions for classification [27]. Using correctly classified test images, we progressively removed top-ranked patches from highlighted heatmap regions and measured the resulting drop in class confidence. For DR detection, the sparse BagNet-Transformer performed best, while standard ViTs performed worst followed by the ResNet-Transformer (Fig. 3e). For AMD severity classification, the sparse ResNet-Transformer outperformed the sparse BagNet-Transformer (Fig. 3f). This may have been due to larger lesion sizes in AMD, favoring CNNs with larger receptive fields. Notably, this aligned with the classification performance, where the ResNet backbone also performed well.
3.5 Our model enhances interpretability for multi-class tasks
Figure 4: Examples of multi-class explanations. Class-specific heatmaps for a Severe DR example from the Kaggle dataset. The first row displays the attention map and corresponding heatmaps from the dense hybrid model with the BagNet backbone, while the second row shows the attention map and heatmaps from its sparse version.
Finally, we visualized class-specific explanations for the dense and sparse BagNet-Transformer. For DR prediction on the Kaggle dataset, both models correctly classified the example (Fig. 4). Heatmaps and class probability distributions were generated in a single forward pass, with the sparse model producing more focused and localized explanations aligned with the predicted class. In other classes, the sparse model showed almost no positive activations, unlike the dense model, which presented a mix of positive and negative evidence. Interestingly, we observed a strong correlation between attention maps and predicted evidence maps, particularly in the sparse model. This suggests that the model effectively captures long-range dependencies in an interpretable way.
4 Discussion and Conclusion
Here, we introduced the first inherently interpretable hybrid CNN-transformer architecture for medical image classification and applied it to AMD severity classification and DR detection from retinal fundus images. Our model was evaluated on two CNN backbones – ResNet, which already captured global spatial relationships, and BagNet, which relied on aggregation of small local features. The latter was particularly interesting as the SA mechanism could help to address the limited receptive field size of the BagNet. Our transformer module employed dual-resolution convolutional self-attention to capture both global and fine-grained features while preserving strong local inductive biases. Unlike standard models with FCL classifiers, our model used an explicit class evidence layer, producing spatial feature maps that serve as class-evidence heatmaps, enabling direct explanation without post-hoc methods.
Interestingly, and in contrast to sparse BagNet models [14], the interpretability vs. accuracy trade-off was relatively small – all evaluated models performed fairly close to each other with high balanced accuracy and $\kappa$, while the sparse BagNet-Transformer clearly produced the best explanations for DR detection performance, while sparse ResNet-Transformer produced the best explanations for AMD severity classification.
Preliminary experiments revealed that multi-head self-attention increased training time without improving classification, and multi-scale resolution had a limited impact. Following [14], we observed that higher sparsity often led to failures in detecting late stages, especially in DR detection (Fig. 4h), likely due to their underrepresentation in the training set. However, our hybrid CNN-Transformer architecture mitigated this issue, demonstrating its effectiveness in low-data settings. Overall, our work underscores hybrid CNN-Transformer models as a strong alternative to post-hoc ViT explanations, particularly for medical imaging.
{credits}
4.0.1 Acknowledgements
This project was supported by the Hertie Foundation, the German Science Foundation (Excellence Cluster EXC 2064 “Machine Learning—New Perspectives for Science”, project number 390727645). The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting KD. PB is a member of the Else-Kröner-Kolleg “ClinBrAIn”.
[…]
The authors declare no competing interests.